OpenStudy (anonymous):

I need help to express these as a+bi... 1) (4(cos(7pi/9)+i sin(7pi/9)))^3 2) (1/2(cos(72degrees)+sin(72degrees)))^5 3) complex 5th roots of 5-5root(3)i 4) all seventh roots of unity

OpenStudy (anonymous):

for the first one, take \(4^3\) and get \(64\) then multiply each angle by 3 and compute \[64\left( \cos(\frac{7\pi}{3})+i\sin(\frac{7\pi}{3})\right)\]

OpenStudy (anonymous):

Oh ok and I would do the same thing for #2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks so much! but I still don't understand the last two

OpenStudy (anonymous):

complex 7th roots of unity divide the circle up in to 7 equal parts, one of which is at 1 since 1 is a seventh root of 1

OpenStudy (anonymous):

this is very easily expressed in trig form as \[\cos(\frac{2\pi}{7})+i\sin(\frac{2\pi}{7})\] and so on

OpenStudy (anonymous):

by "so on" i mean \[\cos(\frac{4\pi}{7})+i\sin(\frac{4\pi}{7})\] etc etc

OpenStudy (anonymous):

to write in the form of \(a+bi\) is i guess a calculator exercise, since i have no idea what \(\cos(\frac{2\pi}{7})\) is

OpenStudy (anonymous):

I know you're supposed to multiply integers 0 - 6 by something... is that something 2pi?

OpenStudy (anonymous):

as for #3 you need to write \[5-5\sqrt3 i\] it polar form first

OpenStudy (anonymous):

yeah i guess it is \(2\pi\) really what you are doing is dividing the unit circle (in the complex plane) in to seven equal parts

OpenStudy (anonymous):



OpenStudy (anonymous):

it is a little hard to draw 7 equal parts, but it should be clear that all the angles will be \[\frac{2\pi}{7}, \frac{4\pi}{7},\frac{6\pi}{7},\frac{8\pi}{7},\frac{10\pi}{7},\frac{12\pi}{7}\] and of course \(0\)

OpenStudy (anonymous):

oh ok thanks so much! :)

OpenStudy (anonymous):

as for #3 you need to write \[5-5\sqrt3 i\] is trig form do you know how to do that?

OpenStudy (anonymous):

I got r=10 and theta=pi/3

OpenStudy (anonymous):

\[r=\sqrt{5^2+(5\sqrt{3})^2}\] yeah 10 looks good

OpenStudy (anonymous):

\(\frac{\pi}{3}\) is not right however

OpenStudy (anonymous):

negative pi/3?

OpenStudy (anonymous):

yeah you can use that one

OpenStudy (anonymous):



OpenStudy (anonymous):

yay thanks

OpenStudy (anonymous):

yw